OpenStudy (anonymous):

Four bells begin to toll together and then each one at intervals of 6 s, 7 s, 8 s and 9 s respectively. The number of times they will toll together in the next 2 hr is: Op 1: 14 times Op 2: 15 times Op 3: 13 times Op 4: 11 times

OpenStudy (anonymous):

take lcm of the seconds

OpenStudy (anonymous):

least common multiple is \[2^3\times 3^2\times 7=504\] so they chime together every \(504\) seconds

OpenStudy (anonymous):

there are \(2\times 60\times 60=7200\) seconds in two hours divide

OpenStudy (kc_kennylau):

wow i'm imagining the situation, so funny xD

OpenStudy (anonymous):

ya got it . thank you.